OpenStudy (anonymous):

The perimeter of a triangle is 30 cm. If the circumference of its incircle is 44 cm then find the area of the triangle

OpenStudy (anonymous):

Some help please.....

OpenStudy (gorv):



OpenStudy (gorv):

2*pi*r=44 pi=22/7 we knw 2*22/7*r=44 44*r/7=44 r=7

OpenStudy (anonymous):

I've already done that but how do you find the sides of the triangle?

OpenStudy (gorv):

perimeter =sum of sides of triangle

OpenStudy (gorv):

30=a+b+c

OpenStudy (gorv):

is anything given about triabgle like equilateral or isosceles or anything ???

OpenStudy (anonymous):

No....... That's the problem

OpenStudy (gorv):

area=r*s r=inradius=7 s=(sum of side of triangle)/2 s=(a+b+c)/2=30/2=15 area=15*7=105